OpenStudy (anonymous):

Give the range of x - 2y = 0 for the domain {-1, 0, 4}.

OpenStudy (anonymous):

D= {-1, 0, 4}. ( I think the range is y) x - 2y=0 (-1)-2y=0 -2y=1 y=-1/2 <---- x - 2y=0 0-2y=0 -2y=0 y=0 <----- x-2y=0 4-2y=0 -2y=-4 y=2 <-----